OpenStudy (anonymous):

solve the inequality -(5/6)

OpenStudy (anonymous):

add 2 to both all three parts \[-\frac{5}{6}+2<x<\frac{1}{8}+2\]

OpenStudy (anonymous):

i guess you should do the arithmetic to make your teacher happy. get \[\frac{7}{6}<x<\frac{17}{2}\]